Derive the relation between magnetic field intensity
 

Grade:12

1 Answers

Avishek Bhagat
3 years ago
Dear Student,

Magnetic intensity and magnetization are related by the relation

M=χH

Also μ=μo
(1+χ)=μr
μo=μr
⟹μr=1+χ